OpenStudy (anonymous):

PLEASE HELP! A survey found that women's heights are normally with mean of 6.28 in. and standard deviation of 2.4 in. The survey also found that men's heights are normally distributed with a mean 68.3 in. and a standard deviation of 2.9 in. Most of the live characters at an amusement park have height requirements with a minimum of 4 ft 8 in. and a maximum of 6 ft 2 in. A. Find the percentage of meeting the height requirement. B. Find the percentage of men meeting the height requirement. C. The percentage of men meeting the height requirement is=

OpenStudy (anonymous):

Since the height requirements are in ft and in, and the mean and standard deviation for both women's height and men's height are in inches, we need to convert the height requirements to inches first. Remember that 1 ft = 12 inches. For example if we have 3 ft and 5 in it will be: 3 * 12 = 36 inches So, 3 ft and 5 inches = 36 + 5 = 41 inches. Is that part clear? :)

OpenStudy (anonymous):

Then, try applying the formula for the z-score:\[z = \frac{x-mean}{sd}\]

OpenStudy (anonymous):

Try finding the z-score for each of the height requirements (which must be in inches first),and use the z-table.

OpenStudy (anonymous):